# OccFiner: Offboard Occupancy Refinement with Hybrid Propagation

### v1

@@ 4.3 Ablation Studies
Table 5: Ablation experiments. Settings used in our final framework are underlined. See Sec. 4.3 for details.
@@
The efficacy of each module in the offboard OccFiner framework is quantified through ablation studies detailed in Tab. 5. Using VoxFormer [27] as the onboard component, we evaluate mIoU accuracy across different distances.
@@
Multi-to-Multi Local Propagation Network. 1) DualFlow4D Transformer. We validate the critical role of the 4D Transformer for local SSC feature propagation, evidenced by a $0.56\%$ drop in mIoU upon its removal. 2) Temporal Embedding, however, did not positively impact accuracy. 3) Spatial Embedding. The result highlights the importance of spatial embedding and the introduction of a long-distance reference frame in the multi-to-multi paradigm, improving accuracy by $0.12\%$. 4) Reference Frame. OccFiner introduces a long-distance reference frame in addition to the local clip, which further improves the accuracy by $0.22\%$. It proves the positive effect of long-term geometric and semantic cues on error compensation. 5) Overall, OccFiner’s first stage, while adding few model parameters ($4.51$M vs. $551$M), effectively compensates for onboard model errors and propagates high-confidence features, boosting mIoU from $13.33\%$ to $14.52\%$, and uniformly improves mIoU across various distances. This is critical for the next global propagation stage.
@@
Region-centric Global Propagation. 1) Average Fusion. Average fusion applied as a baseline to the onboard model yielded only a modest mIoU improvement at ‘far’, indicating limitations at near and middle distances with $0.95\%$ and $0.22\%$ mIoU drop, respectively. 2) Learning Refine. Removing the local propagation network resulted in a significant drop in mIoU by $2.04\%$, highlighting the critical role of first-stage error compensation in enhancing global propagation accuracy across various distances. 3) Sensor Bias. the frustum weighting and distance-aware weighting adjustments are shown to be vital, with their removal leading to noticeable accuracy drops. This underscores the importance of considering the camera field of view and depth uncertainties in long-distance measurements. 4) Temporal Input. The optimal number of frames for global propagation is identified around $(-25,25)$, with more frames offering no additional benefit. 5) Overall, learning-free global propagation significantly improves the accuracy of long-distance prediction. Compared with ‘OccFiner Stage 1 + Vanilla Registration’, OccFiner’s region-centric global propagation further improving the accuracy across different distances by $0.83\%$, $1.10\%$, and $1.98\%$.
5 Conclusions and Discussion
@@
In addressing the challenges of inferior performance and data closure in vision-based SSC, we introduce OccFiner, the first offboard SSC setup to solve the unreliability of the onboard model. By removing computational constraints, OccFiner can reason all frames together and build multi-view consistent SSC predictions. Concretely, OccFiner leverages a hybrid propagation strategy. In the multi-to-multi local propagation stage, OccFiner focuses on the onboard model error compensation and implicit semantic and geometric cues propagation. In the region-centric global propagation stage, OccFiner considers sensor bias and manages long-term information aggregation. Our experimental findings validate OccFiner’s versatility across various onboard models and its capability to significantly elevate SSC quality, setting new benchmarks in vision- and LiDAR-based algorithms on SemanticKITTI. We hope that our framework can become a potent augmentor for onboard SSC algorithms and inspire future offboard research.
@@
Limitations. Our offboard framework, OccFiner, achieves notable results but requires individual retraining for each onboard model. This process is tailored to address the specific errors of each onboard model, with its effectiveness partly dependent on the scene’s characteristics. Additionally, there is substantial scope for improving offboard SSC accuracy in comparison to the ground truth.
@@
Broader impact, Ethics. While OccFiner improves 3D scene understanding, potential inaccuracies in its predictions could have serious implications, especially in critical applications like autonomous driving. It is crucial that such technologies be complemented by additional safety measures to mitigate risks associated with these inevitable errors.

### v4

@@ IV-C Auto-labeling Experiments
TABLE VIII: Ablation experiments. Settings used in our final framework are underlined. See Sec. IV-E for details.
@@
IV-D City-level SSC Map Comparisons
Previous works usually focus on local 3D semantic occupancy, i.e., onboard perception centered on a single vehicle. However, for multi-vehicle collaborative sensing, task scheduling, and traffic control, establishing a city-level SSC map is of great significance. Therefore, we propose to fuse local SSC predictions using relative poses to further explore the effectiveness of OccFiner in establishing visually derived city-level SSC results. In the global map, we only consider static rigid targets of the scene, since dynamic targets are not among the usual properties of maps.
Quantitative Comparisons: As shown in Table VII, we benchmark the mainstream LiDAR- and camera-based SSC methods, and explore OccFiner + VoxFormer to establish an offboard city-level SSC Map. Experimental results show that the accuracy of the LiDAR-based SSC method is consistently ahead of the camera-based method in city-level SSC maps. It is worth noting that OccFiner further improves the accuracy of both LiDAR- and camera-based city-level SSC maps. Compared with the onboard component [4], the geometric IoU of OccFiner (Vox) increased by $0.93$%, and the semantic mIoU improved by $1.51$%. It even exceeds the LiDAR-based SSCNet-full [18] by $0.76$% in geometric accuracy and surpasses the LiDAR-based SSCNet [18] by $0.83$% in semantic accuracy. This once again proves that the proposed OccFiner framework can effectively unleash the potential of purely visual SSC perception solutions.
Qualitative Comparisons: We visualize the validation sequence of SemanticKITTI [16] to evaluate the quality of city-level Semantic Scene Completion (SSC) maps. The spatial resolution of the original single-frame local occupancy is $256\times 256\times 32$. After splicing and voting using relative pose registration, the spatial resolution of the global occupancy map is $2205\times 4296\times 261$, covering a spatial range of $441\times 859.2\times 52.2~{}m^{3}$. As illustrated in Fig. 6, directly using the outputs of the visual SSC algorithm (VoxFormer [4]) for global registration, despite their large onboard errors, leads to artifacts on key road surfaces. The implementation of OccFiner significantly reduces these errors, thereby enhancing the quality of the city-level SSC maps. Additionally, Fig. 7 presents the global map visualized from a Bird’s-Eye View (BEV) perspective. The VoxFormer [4] results exhibit notable radial artifacts, attributed to the limited viewing angle of the camera’s view frustum. In contrast, OccFiner’s region-centric global propagation successfully compensates for the camera’s measurement limitations, effectively eliminating these radial artifacts.
IV-E Ablation Studies
@@
The efficacy of each module in the offboard OccFiner framework is quantified through ablation studies detailed in Tab. VIII. Using VoxFormer [4] as the onboard component, we evaluate mIoU accuracy across different distances.
@@
Multi-to-Multi Local Propagation Network: 1) DualFlow4D Transformer. We validate the critical role of the 4D Transformer for local SSC feature propagation, evidenced by a $0.56\%$ drop in mIoU upon its removal. 2) Temporal Embedding, however, did not positively impact accuracy. 3) Spatial Embedding. The result highlights the importance of spatial embedding and the introduction of a long-distance reference frame in the multi-to-multi paradigm, improving accuracy by $0.12\%$. 4) Reference Frame. OccFiner introduces a long-distance reference frame in addition to the local clip, which further improves the accuracy by $0.22\%$. It proves the positive effect of long-term geometric and semantic cues on error compensation. 5) Overall, OccFiner’s first stage, while adding few model parameters ($4.51$M vs. $551$M), effectively compensates for onboard model errors and propagates high-confidence features, boosting mIoU from $13.33\%$ to $14.52\%$, and uniformly improves mIoU across various distances. This is critical for the next global propagation stage. In addition, the local propagation network can also be directly used for LiDAR-based SSC prediction. We add the SSC accuracy of the base model under various time window settings. As shown in Tab. IX, the accuracy consistently improves with an increase in the time window, while keeping the model structure unchanged. This further underscores the effectiveness and rational design of the base model structure.
TABLE IX: Performance of directly using multi-to-multi propagation network with different time window.
Dataset Input Time Window mIoU
SemanticKITTI LiDAR Scans 1 17.99
2 19.10
4 19.98
6 20.68
@@
Region-centric Global Propagation: 1) Average Fusion. Average fusion applied as a baseline to the onboard model yielded only a modest mIoU improvement at ‘far’, indicating limitations at near and middle distances with $0.95\%$ and $0.22\%$ mIoU drop, respectively. 2) Learning Refine. Removing the local propagation network resulted in a significant drop in mIoU by $2.04\%$, highlighting the critical role of first-stage error compensation in enhancing global propagation accuracy across various distances. 3) Sensor Bias. the frustum weighting and distance-aware weighting adjustments are shown to be vital, with their removal leading to noticeable accuracy drops. This underscores the importance of considering the camera field of view and depth uncertainties in long-distance measurements. 4) Temporal Input. The optimal number of frames for global propagation is identified around $(-25,25)$, with more frames offering no additional benefit. 5) Overall, learning-free global propagation significantly improves the accuracy of long-distance prediction. Compared with ‘OccFiner Stage 1 + Vanilla Registration’, OccFiner’s region-centric global propagation further improving the accuracy across different distances by $0.83\%$, $1.10\%$, and $1.98\%$.
The proposed Occfiner, as the first offboard occupancy approach, integrating both local and global propagation techniques, not only demonstrates significant improvements in predictive accuracy but also establishes a robust framework for achieving data loop-closure in intelligent transportation systems. By addressing onboard model deficiencies and enhancing error compensation through offboard processing, our methodology lays a solid foundation for future advancements in self-evolving 3D scene understanding technology, promoting safer and more reliable navigation.
V Conclusions and Discussion
@@
In addressing the challenges of inferior performance and data loop-closure in vision-based Semantic Scene Completion (SSC), we introduce OccFiner, the first offboard SSC setup designed to enhance the reliability of onboard models. By removing computational constraints, OccFiner reasons with all frames together to build multi-view consistent SSC predictions. It employs a hybrid propagation strategy: compensating for onboard model errors and propagating semantic and geometric cues locally, while managing sensor bias and long-term information aggregation globally. Our experiments validate OccFiner’s versatility across various onboard models and its ability to significantly elevate SSC quality, setting new benchmarks on SemanticKITTI for both vision- and LiDAR-based algorithms. Integrating data into offboard occupancy systems successfully addresses model deficiencies, achieves data loop-closure, and enables system self-evolution. This approach also solves the distribution shift between training and deployment, allowing for scalable, cost-effective improvements in the accurate reconstruction of 3D scenes. We hope that our framework will augment onboard SSC algorithms and inspire future offboard research, advancing autonomous driving technology by reducing manual annotation efforts and improving large-scale 3D scene understanding.
Implications: This study advances autonomous driving perception by demonstrating three key insights: (1) Vision-based systems with offboard processing (OccFiner) achieve LiDAR-comparable 3D scene understanding, challenging LiDAR’s necessity for geometric reconstruction. (2) Camera-derived semantic scene completion enables scalable vision-centric systems through automated pseudo-labeling, reducing manual annotation dependence. (3) Offboard processing enhances scene representation completeness and consistency, critical for safety decisions and map automation.
@@
Limitations: Our offboard framework, OccFiner, achieves notable results but requires individual retraining for each onboard model. This process is tailored to address the specific errors of each onboard model, with its effectiveness partly dependent on the scene’s characteristics. Additionally, there is substantial scope for improving offboard SSC accuracy in comparison to the ground truth. Furthermore, it must be acknowledged that accurate perception, especially for fine-grained geometry and semantics using only cameras, remains challenging at longer distances (e.g., towards the $51.2m$ evaluation boundary) due to inherent limitations in resolution and depth estimation, contributing to lower performance in far-range evaluations.
Recommendations for Future Work: Our framework demonstrates that offboard occupancy systems enable data loop-closure and mitigate distribution shifts, achieving cost-effective scalability in 3D reconstruction while reducing manual annotation burdens. Future work should optimize OccFiner’s cross-model generalization, bridge residual accuracy gaps through advanced learning, and integrate refined outputs into decision systems. Ultimately, unified multitask offboard architectures could revolutionize perception.
@@
Broader impact, Ethics: While OccFiner improves 3D scene understanding, potential inaccuracies in its predictions could have serious implications, especially in critical applications like autonomous driving. It is crucial that such technologies be complemented by additional safety measures to mitigate risks associated with these inevitable errors.